Johns budget line relating to good x and good y has intercept of 50 units of good x and 20 units of food y. If the price of good x is $12 what is johns income?
1
Expert's answer
2020-05-25T09:07:45-0400

The correct answer is:

If a budget line relating good X and good Y has intercept Qx = 50 unit of good X and Qy = 20 units of good Y and Px = $12, then the John's income is:

I = Px*Qx + Py*Qy

When the budget line has the intercept Qx = 50 units and Px= $12, then I = 12*50 + py*0= 600

so the John's income is $600.